### v2

# Modeling the amplification of epidemic spread by misinformed populations

Abstract
Understanding how misinformation affects the spread of disease is crucial for public health, especially given recent research indicating that misinformation can increase vaccine hesitancy and discourage vaccine uptake. However, it is difficult to investigate the interaction between misinformation and epidemic outcomes due to the dearth of data-informed holistic epidemic models. Here, we propose an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data. Our model allows us to simulate and estimate various scenarios to understand the impact of misinformation on epidemic spreading. Using this model, we estimate that misinformation could have led to 47 million additional COVID-19 infections in the U.S. in a worst-case scenario.
Social factors, such as information sharing, play a crucial role in shaping the dynamics and epidemiology of infectious diseases12, 5. For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures32, 4. A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic38, or by social media and the anti-vaccination movement1, 24, 10, 13.
In particular, misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance43. Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States40, 37. Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments31. This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity16, 45. Proper management of epidemic crises in the modern age thus requires the understanding of the complex relationship between the spread of (mis)information through online social networks and the spread of disease through physical contact networks (Fig. 1).
Figure 1: The spread of misinformation affects the transmission of disease. (A) Schematic illustration of the misinformation and contact networks. Online social networks foster misinformation dissemination while physical contact networks, such as those that connect co-workers in an office or pupils in a school, facilitate disease transmission. Dotted links indicate that the same people participate in both networks, which have different topologies; e.g., the information network tends to have stronger political homophily while the contact network tends to have stronger geographic homophily. We focus on the impact of misinformation spread on disease transmission (downward arrow), while the opposite effect (upward arrow, e.g., individuals ceasing to share misinformation due to illness) falls outside the scope of this investigation. (B) A contact network based on 0.01% county population samples. Nodes are sized based on degree (number of contacts). In a scenario with limited spread of misinformation (black nodes in C), the simulations of disease spread leads to a number of infected individuals (red nodes in D). In a scenario where the misinformation spreads more widely (purple nodes in E), more individuals get infected (yellow nodes in F).
While previous agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways48, 34, 39, 9, there is a growing need to integrate real-world data to strengthen the connections between simulation results and real-world outcomes. Social media can provide valuable data sources for this purpose6, 49.
We address this challenge by proposing a data-informed epidemic model that takes both the distribution of misinformed individuals and a physical mobility network into account. Using this data, we augment the Susceptible-Infected-Recovered (SIR) model to account for a subpopulation of “misinformed” individuals, who do not heed expert public health guidance. We explore how this group can affect the larger, ordinary population using both a mean-field approximation, which assumes all individuals have an equal chance of interacting, and a multi-level agent-based simulation based on two large, data-informed network: a social network where misinformation spreads and a contact network where the disease can propagate. A contact network of approximately 20 million nodes is constructed by leveraging large-scale Twitter data, county-level voting records, and cellphone mobility data. We incorporate theoretically extreme values of the parameter responsible for the epidemic transmission to evaluate best- and worst-case scenarios about the impact of misinformed individuals on the spread of disease and obtain quantitative bounds on the harm caused by misinformation.
Results
In our model, as in the standard SIR modeling framework2, a parameter $\beta$ describes the average number of infected individuals generated by an infected individual in a time unit. Infected individuals recover with rate $\gamma$. We extend this epidemic model to account for misinformed and ordinary subpopulations. We utilize six compartments to capture the three stages of SIR for each subpopulation. The fractions of individuals in the susceptible, infected, and recovered compartments are denoted by $S_{O}$, $I_{O}$, and $R_{O}$ respectively for the ordinary group; and by $S_{M}$, $I_{M}$, and $R_{M}$ for the misinformed group.
Ordinary individuals are considered to be well-informed about public health guidelines, while misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow recommended behaviors, such as social distancing, mask-wearing, and vaccination, thereby increasing the risk of infection for themselves and others19. We therefore define two distinct parameters $\beta_{O}$ and $\beta_{M}$ for ordinary and misinformed susceptibles, respectively. We model the risky behavior of the misinformed group by setting $\beta_{M}\geq\beta_{O}$. The recovery rate $\gamma$ is identical for both subpopulations. We refer to this as the Susceptible Misinformed Infected Recovered (SMIR) model.
Mean-field simulations
The mean-field approach assumes that an infected individual can transmit disease to any susceptible individual. With this assumption, we derive the model’s mean-field system of equations (see Methods) and explore parameter values for which epidemic spread occurs. We denote the proportions of ordinary and misinformed individuals as $\mu$ and $1-\mu$, respectively. An initial proportion of the population $\epsilon$ is infected, split evenly between the ordinary and misinformed groups. Thus, we initialize the compartments with $I_{O}=\frac{\epsilon}{2}$, $I_{M}=\frac{\epsilon}{2}$, $S_{O}=\mu-\frac{\epsilon}{2}$, $S_{M}=1-\mu-\frac{\epsilon}{2}$, and $R_{O}=R_{M}=0$.
The effects of varying the transmission and recovery parameters are predictable (see Supplementary Information): lower $\beta$ and higher $\gamma$ “flatten the curve” and reduce the negative outcomes of an infection. Based on these explorations we set $\beta_{O}=0.3$ and $\gamma=0.2$ such that the basic reproduction number $R_{0}=\beta/\gamma>1$ to ensure epidemic spread. Next we focus on the effects of varying the misinformed subpopulation’s level of risk and the degree of homophily within the network.
Figure 2: Increasing $\lambda=\beta_{M}/\beta_{O}$ accelerates and amplifies the infection. We use $\beta_{O}=0.3$, $\gamma=0.2$, and $\mu=0.5$. (A) Proportion of the population infected on each day. (B) Extra proportion of the total population that is infected as a function of $\lambda$.
To explore the effects of risky behaviors by misinformed individuals, let us assume two equally-sized subpopulations ($\mu=1/2$) and introduce the scaling factor $\lambda=\beta_{M}/\beta_{O}\geq 1$. Fig. 2 illustrates the increasing negative impact of the misinformed subpopulation on the disease-spreading dynamics as $\lambda$ becomes larger. The social cost associated with the more risky behaviors by the misinformed group is passed on to the whole network. For example, when $\lambda=3.0$, peak infection for the entire population is reached 38 days earlier than in the $\lambda=1$ case (day 22 vs. 60; Fig. 2A), leading to an additional 29.2% of the population becoming infected (Fig. 2B).
Let us explore the effect of homophily among the ordinary and misinformed subpopulation networks. Homophily means that infected individuals are more likely to interact with (and infect) susceptibles from the same subpopulation (ordinary or misinformed) than the other group. The degree of homophily is modeled by a parameter $\alpha$. When $\alpha=0.5$, individuals are equally likely to interact within and across groups (no homophily), whereas $\alpha=1$ is the case when homophily is strongest and the subpopulations do not interact with each other (see Methods for details).
Figure 3: Homophily in the contact network worsens the infection among misinformed individuals, especially for lower transmission rates. The combined effects of transmission and homophily parameters, $\beta_{O}$ and $\alpha$, are examined with the mean-field approximation when $\lambda=3$, $\gamma=0.2$, and $\mu=0.5$. We plot the proportions of infected individuals in (A) the ordinary population, (B) the misinformed population, and (C) the overall population. The maximum proportion of the overall population infected for each $\beta_{O}$ is marked with a black dot. When the transmission rate is sufficiently high, homophily benefits the entire population but harms the misinformed group.
Fig. 3 illustrates the effects of homophily ($0.5\leq\alpha\leq 1$) for different levels of ordinary transmission ($0.1\leq\beta_{O}\leq 0.4$). For less infectious disease (low $\beta_{O}$), increasing homophily significantly harms the misinformed group (Fig. 3B): the infection remains confined within this group. There is no discernible effect on the ordinary population as long as the two groups interact; when they do not ($\alpha=1$), we observe a sharp reduction in infections (Fig. 3A). For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig. 3C 35. Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig. 3B and C). As $\beta_{O}$ increases further, while nearly the entire misinformed population becomes infected regardless of $\alpha$ (Fig. 3B), high homophily shields the full population (Fig. 3C): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals. In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 44, 47.
Agent-based simulations
To examine the influence of misinformation within a more realistic epidemic framework, we utilize a multi-level model. Our approach combines an empirically derived information network with a contact network calibrated with real-world data, as illustrated in Fig. 4. Information diffusion is modeled by leveraging a large set of users of a popular social media platform and epidemic simulations are subsequently conducted on contact networks populated with misinformed individuals.
Figure 4: An idealized example of our multi-level modeling framework. (A) Spread of misinformation through an information network (dashed lines). Colors represent ideological homophily. Nodes with bold borders are misinformed about the epidemic. The misinformation spreads through a complex contagion (linear threshold) model; two scenarios show that a lower threshold $\phi$ leads to more misinformed nodes. (B) Construction of the contact network (solid lines) for counties with sufficient information diffusion data (in black) to provide reasonable estimates about the fraction of misinformed individuals. Note that these counties account for 63.52% of U.S. voters. Each location’s population size and ideological mix are based on empirical data, and misinformed individuals are based on the information diffusion model. Links among individuals within and between locations are based on empirical mobility data. (C) The infection spreads through the contact network (black nodes), according to the SMIR model.
We leverage a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines21. From approximately nine months of this data (Jan. 4–Sep. 30, 2021), we geolocate over 2 million U.S. users who shared almost 26 million tweets and focus on accounts in 341 U.S. counties containing more than 200 Twitter users each. We also infer an account’s political alignment and whether they shared any likely misinformation (see Methods). Twitter is not representative of the U.S. population, and people also access information in other ways, such as traditional media and word of mouth. However, this empirical social network serves as a data-informed proxy for a large, realistic network through which people share information about the disease.
With this data, we build a directed and weighted information diffusion network, in which an edge $(i\rightarrow j,w)$ indicates that $j$ retweeted $i$ $w$ times. There are various ways to model the infodemic18. We simulate the spread of misinformation on this network, as illustrated in Fig. 4A. Accounts that share or reshare posts containing misinformation are considered misinformed. These accounts function as the initial seeds from which misinformation proliferates to the wider network. Not all users actively participate in content sharing, with only about half of U.S. Twitter users, for instance, engaging in this activity36. But even without active sharing, exposure to misinformation can still influence individual behavior31.
To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process25. This is a simple way to model complex contagion15, 51, 33, according to which individuals require multiple exposures to misinformation before they become misinformed themselves. Let a linear threshold $\phi$ represent the minimum number of misinformed friends needed for an ordinary node to become misinformed. If the total number of misinformed friends of $i$ is greater than or equal to $\phi$, $i$ is marked as misinformed ($M$). The remaining nodes are marked as ordinary susceptibles ($O$). We can interpret $\phi$ as a measure of “resilience” to misinformation; as it increases, individuals require more exposure to misinformation to be converted to the misinformed group. Note that since we explore the full range of $\phi$ values, the following results are unaffected whether the threshold is defined based on the number of users or the number of retweets.
Fig. 5A shows how $\phi$ influences the number of misinformed individuals within the retweet network. With strong resilience ($\phi>10$), exposure to misinformation does not have much effect and few nodes are converted to the misinformed group. Conversely, when resilience to misinformation is very low (as in the simple contagion case $\phi=1$), all nodes exposed to a misinformation-containing post are converted to the misinformed group. Through this process, empirically observed misinformation-sharing behavior leads to information networks with misinformed subpopulations of varying sizes based on different $\phi$ values.
Figure 5: More misinformed individuals lead to a larger portion of the network becoming infected. Decreasing the resilience $\phi$ (A) increases the size of the misinformed subpopulation, leading to (B) faster infection spreading and (C) a greater cumulative number of infections. In panels (B, C), lines and corresponding shaded regions represent the mean and standard deviation across simulations, respectively.
We generate contact networks for different thresholds ($1\leq\phi\leq 20$) to compare the impact of misinformed subpopulations of different sizes. Given a threshold $\phi$ and the corresponding information network, we aim to construct a physical contact network containing empirically calibrated misinformed subpopulations (Fig. 4B). The process begins by selecting a sample of individuals from each county within the information network. As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic27, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S. presidential election. For each county, we add the sampled nodes to the physical network marked as misinformed ($M$) or ordinary susceptible ($O$), based on their label within the retweet network. Sampling with replacement allows us to select individuals such that the overall proportions of Republicans and Democrats match the voting records. A 10% sample leads to $N\approx 20$ million nodes. A network based on a much smaller sample is illustrated in Fig. 1B. This process captures empirical measurements of the ideological split, relative population size, and quantity of misinformed individuals in each county. We add contact network edges by leveraging cell phone mobility data that provides the probability of an individual traveling within and between counties. See Methods for details.
Disease-spreading dynamics on a contact network are simulated using an agent-based version of SMIR (Fig. 4C). Two critical parameters affecting these dynamics are the density of the contact network, captured by its average degree $\bar{k}$, and the transmission probability $p$ (note that $p\bar{k}=\beta$). Mitigation measures such as social distancing decrease $\bar{k}$, while those such as masking and vaccination decrease $p$. A simple way to model the combined effects of misinformation on these behaviors is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic30, and explore a wide range of transmission probabilities to capture worst- and best-case scenarios. To this end, we model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals. In contrast, we model the adoption of several mitigation measures by selecting an extremely small value $p_{O}=0.01$ for ordinary individuals. These values illustrate two theoretically extreme scenarios. The former portrays a realistic number of interactions during non-pandemic times, accompanied by high transmission rates due to the absence of preventive measures, such as social distancing, mask-wearing, or vaccinations. The latter demonstrates decreased daily interactions and reduced transmission rates resulting from the implementation of these preventive measures. Using the empirically calibrated contact networks in conjunction with these extreme parameters, the simulation approach allows us to bound the best- and worst-case scenarios in a data-informed manner (see Methods for more information).
The worst-case effect of the misinformed subpopulation size on the daily incidence of infection (illustrated in Fig. 1C-F on a small network) is quantified in Fig. 5B on a large network (10% sample). A heavily misinformed population ($\phi=1$) corresponds to an additional 9% of the population being infected at peak time (a six-fold increase) compared to a population following expert guidance in the best-case scenario ($\phi=20$). The peak also occurs approximately two weeks earlier. The cumulative effect of a heavily misinformed population is also significant, with an additional 14% of the population infected over the course of the epidemic — a 32% relative increase (Fig. 5C). We repeated this analysis for different sample sizes and found that the main results are robust (see Supplementary Information).
Discussion
This work presents large-scale simulations based on an epidemic model that accounts for a subpopulation of misinformed individuals, informed by real-world data. Such an approach answers calls from the literature to anchor models in empirical data6, 49, allowing us to explore realistic potential outcomes. While we know that exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal37, our model further links these behaviors by a misinformed population to an accelerated spread of disease.
By deriving the mean-field system of equations for the novel SMIR model and exploring its dynamics, we demonstrate how the risky behaviors of misinformed individuals can adversely impact those following public health guidelines, worsening outcomes for the entire population. Additionally, increasing homophily can benefit the overall population by protecting ordinary citizens; however, it may also lead to higher infection rates within the misinformed subpopulation.
Agent-based simulations of the SMIR model let us study the epidemic on empirically calibrated contact networks, which provide more realistic scenarios compared to the mean-field approach. The model estimates that, in the worst-case scenario, a heavily misinformed population can amplify the peak of the infection by a factor of six and accelerate it by two weeks. This would result in an additional 14% of the population becoming infected — nearly 47 million Americans based on recent U.S. Census data50. The corresponding price tag of vaccine misinformation would be over \$143B, using estimated health care costs associated with COVID-19 in the U.S.3.
These results provide an alarming quantitative bound on the harm of exposure to online vaccine misinformation. They should provide public health authorities as well as social media platforms with heightened motivation to curb vaccine misinformation, despite the difficulties posed by social media design11.
We acknowledge certain limitations in our approach. While individual beliefs and behaviors may vary over time, our model simplifies the scenario by dichotomizing individuals into misinformed and ordinary subpopulations and assuming constant transmission rates. We also assume uniform resilience to misinformation for all individuals during the information diffusion process, although this attribute likely differs across individuals. Future directions could involve more sophisticated models to account for these heterogeneities. For instance, cognitive models of misinformation acceptance7 could be incorporated into the simulation with misinformation exposure data collected from social media. Such integration would enable the transition of individuals from ordinary to misinformed susceptible states throughout the simulation, allowing for a simultaneous examination of opinion and disease dynamics. Some theoretical models have already explored similar approaches and obtained results that align with our findings34, 48. Additional extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population.
Methods
SMIR model.
For both the ordinary and misinformed subpopulations, the Susceptible Misinformed Infected Recovered (SMIR) model replicates the standard SIR compartments, denoted as $S_{O}$/$I_{O}$/$R_{O}$ and $S_{M}$/$I_{M}$/$R_{M}$, respectively. While SMIR adopts distinct transmission parameters for the misinformed and ordinary groups ($\beta_{M}$ and $\beta_{O}$, respectively), the mean-field approach assumes that infected individuals from either group ($I_{O}+I_{M}$) can potentially infect anyone in the susceptible populations. The mean-field model is governed by the following system of equations:
$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-\beta_{O}S_{O}(I_{O}+I_{M}),\;\;\frac{dI_{O}}{dt}=\beta_{O}S_{O}(I_{O}+I_{M})-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\\ \frac{dS_{M}}{dt}=-\beta_{M}S_{M}(I_{O}+I_{M}),\;\;\frac{dI_{M}}{dt}=\beta_{M}S_{M}(I_{O}+I_{M})-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (1)
To account for homophily, we modify the term $I_{O}+I_{M}$ in Eq. 1 to account for increased (decreased) contacts within (across) groups, according to the parameter $\alpha\geq 0.5$. When homophily does not play a role ($\alpha=0.5$), there is an equal probability of interacting with either subpopulation’s infected group. We thus obtain:
$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-2\beta_{O}S_{O}(I_{O}\alpha+I_{M}(1-\alpha)),\;\;\frac{dI_{O}}{dt}=2\beta_{O}S_{O}(I_{O}\alpha+I_{M}(1-\alpha))-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\\ \frac{dS_{M}}{dt}=-2\beta_{M}S_{M}(I_{O}(1-\alpha)+I_{M}\alpha),\;\;\frac{dI_{M}}{dt}=2\beta_{M}S_{M}(I_{O}(1-\alpha)+I_{M}\alpha)-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (2)
We initialize the simulation by randomly marking $\frac{\epsilon}{2}$ nodes as infected in both the misinformed and ordinary groups, where $\epsilon=0.001$.
Twitter and derived data.
Twitter posts in the CoVaxxy dataset21 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API). To capture the online discourse surrounding COVID-19 vaccines in English, a comprehensive set of English-language keywords was carefully curated. Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique22 was used to identify co-occurring relevant keywords in December 202021. The resulting list contained almost 80 keywords, available online20. To confirm the relevance of the collected tweets to the topic of vaccines, we examined the coverage obtained by incrementally adding keywords, starting with the most common ones. Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library23 that leverages self-reported location metadata within user profiles (embedded in tweets). As an account’s location may change over time (captured across multiple tweets), we utilize the most recent location. We geolocate 2,047,800 users residing in all 50 U.S. states, who shared a total of 25,806,856 tweets by mapping self-reported locations to U.S. counties. The information network is constructed from accounts in 341 counties that contain more than 200 Twitter users each. Political alignment is estimated using a third-party list of annotated news sources42, 41. It is averaged across all the sources shared by each account. Nodes with an estimated alignment greater (smaller) than zero are considered Republican (Democrat). We infer the political alignment of some additional accounts, who did not share links to news sources, using a label-propagation algorithm17 on the retweet network. If all of a node’s neighbors have political alignment scores, its score is estimated using the weighted average of its neighbors, with weights based on retweets. The process is iterated until each node without a score has at least one neighbor without a score. Misinformation is defined at the source level. Tweets containing links to articles from a list of low-credibility sources compiled by NewsGuard (score below 60) are labeled as spreading misinformation. This approach is common practice and has been validated in the literature37, 26, 8, 29, 46.
Contact network edges.
To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data52, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States. This data has been widely employed to study human mobility patterns during the COVID pandemic. We used the average daily number of individuals moving during 2019, as a reference for business-as-usual mobility, and aggregated all CBGs and POIs at the county level. This aggregation results in a county-by-county matrix $L$, where each element $L_{xy}$ represents the average daily number of individuals in county $x$ moving to county $y$ or vice versa. We then normalized $L_{xy}$ to obtain the average probability of individuals in counties $x$ and $y$ coming into contact, and multiplied by the total number of edges to obtain the expected number of connections between individuals in counties $x$ and $y$: $E_{xy}=\frac{L_{xy}}{\sum_{x^{\prime},y^{\prime}}L_{x^{\prime}y^{\prime}}}\frac{\bar{k}N}{2}$ where the sum is over all county pairs and $\frac{\bar{k}N}{2}$ is the total number of edges. Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model28 used to generate networks with localized communities. For each pair of distinct locations $x$ and $y$, we draw $E_{xy}$ edges between random pairs of nodes in $x$ and $y$. Additionally, we draw $E_{xx}$ edges among random pairs of individuals within the same location $x$, representing homogeneous mixing within each county. At the end of the process, the network has the target average degree $\bar{k}$. We use $\bar{k}=25$ and show how this parameter affects the infections in Supplementary Information.
Simulation details.
Agent-based SMIR simulations are initiated by randomly selecting 100 misinformed nodes and designating them as infected. The disease spreading dynamics are then simulated for 100 steps, which correspond to days. To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period14 ($\gamma=0.2$). Each simulation is repeated ten times, and the average outcome is reported.
Data and code availability.
Code and data are available in a public repository (github.iu.edu/NaN-team/bounding-misinfo-impact-on-disease-spread).
Acknowledgments
We are grateful to Yuan Yuan, Marco Ajelli, Alessio Brina, and Brea Perry for helpful discussions. This work was supported in part by the Swiss National Science Foundation (grant 209250), the National Science Foundation (grants 1927425 and 1927418), the Army Research Office (contract W911NF-21-1-0194), the European Union (NextGenerationEU project PNRR-PE-AI FAIR), the Italian Ministry of Education (PRIN PNRR grant CODE prot. P2022AKRZ9 and PRIN grant DEMON prot. 2022BAXSPY), Knight Foundation, and Craig Newmark Philanthropies.
Ethics
This study, focusing on public data, poses minimal risk to human subjects. Consequently, the Indiana University Institutional Review Board has exempted it from review (protocol number 1102004860). All data collection and analysis adhered to Twitter’s terms of service.
Supplementary information
Appendix 1 Mean-field SMIR model
Let us denote the proportions of misinformed and ordinary individuals as $\mu$ and $1-\mu$, respectively. A proportion $\epsilon=0.001$ of the population is initially infected, split evenly between the ordinary and misinformed groups. Thus, during the initial state, we have initial values for each compartment: $R_{O}=R_{M}=0$, $S_{O}=\mu-\frac{\epsilon}{2}$, $S_{M}=1-\mu-\frac{\epsilon}{2}$, $I_{M}=\frac{\epsilon}{2}$, and $I_{O}=\frac{\epsilon}{2}$.
Figure S1: Reducing the transmission parameter $\beta_{O}$ decreases the severity of the epidemic. We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values, and (c) the total proportion of the population infected as $\beta_{O}$ varies. Here we do not consider the role of misinformation or homophily.
To identify a suitable base value for the transmission rate among ordinary susceptibles, we begin by exploring the scenario with no misinformed individuals ($\mu=1$), setting $\gamma=0.2$ and varying the transmission parameter in the range $0.02\leq\beta_{O}\leq 1$. As is typical of SIR dynamics, Fig. S1 shows that lower $\beta_{O}$ values delay and lower the infection peak — the so-called “flattening of the curve.” Lower $\beta_{O}$ also decreases the total proportion of the population that becomes infected at any point during the epidemic, while higher $\beta_{O}$ values increase this proportion. These dynamics are tied to the basic reproduction number $R_{0}=\beta/\gamma$: the disease spreading dynamics only reach epidemic levels when $R_{0}>1$, such that an infected individual infects more than one other person on average. This happens when $\beta_{O}>0.2$. As $R_{0}$ increases, the infection spreads more quickly, the peak infection day occurs sooner, and the proportion of the population that is ultimately infected increases.
Figure S2: Effects of varying the recovery rate. We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values, and (c) the total proportion of the population infected as a function of the number of days to recover, $\tau$. Here we do not consider the role of misinformation or homophily.
We now explore the effect of the recovery rate, again in the scenario with no misinformation or homophily, by setting $\beta_{O}=0.3$ and varying the recovery period $\tau=1/\gamma$ between 1 and 20 days. Fig. S2 shows that when $\tau<4$, $R_{0}<1$ and the disease does not reach epidemic proportions. At this level, the epidemic takes a long time to reach its peak ($\approx$80 days). Increasing $\tau$ means that individuals remain infected longer, so the population gets infected faster and the peak infection is reached more rapidly. We set $\tau=5$ ($\gamma=0.2$) in the main text to align with quarantine recommendations from the CDC14. With this value and $\beta_{O}=0.3$, the dynamics are at epidemic levels ($R_{0}=1.5$).
Appendix 2 Robustness analyses
To test the robustness of the main results (Fig. 5C) with respect to the sample size used to construct the contact network, all simulations were rerun after generating contact networks based on the different sampling percentages between 0.01% and 10%. Fig. S3 shows the relative increase in the percentage of the population that becomes infected as a function of the linear threshold $\phi$, using the best-case scenario in which the fewest nodes in the network are misinformed as the baseline. We observe a substantial decrease in the effect of misinformation as the sampling size grows to 1%. However, sample sizes above 1% return nearly identical results. We conclude that using a sample size of 10% (as reported in the main text) is sufficient to rule out any size-induced bias.
Figure S3: Relative increase in the mean total population infected as a function of the sampling size utilized in the contact network creation process. The $\phi=20$ scenario, in which the fewest nodes in the network are misinformed, is utilized as the baseline.
We also investigate the impact of the contact network density ($\bar{k}$) on infection dynamics, as illustrated in Fig. S4 for $5\leq\bar{k}\leq 25$. We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions30. As expected, Fig. S4(a) demonstrates that higher $\bar{k}$ leads to increased infections through the population, since the higher contact density provides more opportunities for transmission. But while a larger percentage of the overall population is infected, the relative effect of misinformed individuals decreases. This is because, at higher $\bar{k}$ values, the infected population is already substantial even in the low-misinformation ($\phi=20$) baseline. The combined effect of these two opposing trends, as shown in Fig. S4(b), is that the additional percentage of infected individuals relative to the $\phi=20$ scenario reaches a maximum for some intermediate $\bar{k}$. Fig. S4(b) also shows that, consistent with our primary findings, larger $\phi$ (misinformed resilience) results in a smaller infected population. In our main analysis we focus on $\bar{k}=25$ and model an effective reduction of contacts, resulting for example from social distancing or lockdowns, by decreasing the $p$ parameter.
Figure S4: Effects of average contact network degree $\bar{k}$ on infection dynamics. (a) Infected individuals ($\phi=1$) as a percentage of the overall population and relative to the baseline condition $\phi=20$, in which the fewest nodes in the network are misinformed. (b) Additional percentages of infected population relative to the baseline condition $\phi=20$.

### v3

@@
 However, it is difficult to investigate the interaction between misinformation and epidemic outcomes due to the dearth of data-informed holistic epidemic models.
-Here, we propose an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data.
-Our model allows us to simulate and estimate various scenarios to understand the impact of misinformation on epidemic spreading.
-Using this model, we estimate that misinformation could have led to 47 million additional COVID-19 infections in the U.S.
-in a worst-case scenario.
+Here, we employ an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data.
+The model allows us to simulate and estimate various scenarios to understand the impact of misinformation on epidemic spreading.
+Using this model, we present a worst-case scenario in which a heavily misinformed population would result in an additional 14% of the U.S.
+population becoming infected over the course of the COVID-19 epidemic, compared to a best-case scenario.
 Social factors, such as information sharing, play a crucial role in shaping the dynamics and epidemiology of infectious diseases12, 5.
-For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures32, 4.
-A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic38, or by social media and the anti-vaccination movement1, 24, 10, 13.
-In particular, misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance43.
-Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States40, 37.
-Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments31.
-This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity16, 45.
+For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures34, 4.
+A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic40, or by social media and the anti-vaccination movement1, 26, 10, 13.
+A great deal of work has explored how to model the influence of human behaviour on the spread of infectious diseases53, 25.
+Here we focus on risky behaviors affecting disease transmission that are associated with misinformed individuals.
+Misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance45.
+Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States42, 39.
+Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments33.
+This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity17, 47.
 Proper management of epidemic crises in the modern age thus requires the understanding of the complex relationship between the spread of (mis)information through online social networks and the spread of disease through physical contact networks (Fig.
@@
 In a scenario where the misinformation spreads more widely (purple nodes in E), more individuals get infected (yellow nodes in F).
-While previous agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways48, 34, 39, 9, there is a growing need to integrate real-world data to strengthen the connections between simulation results and real-world outcomes.
-Social media can provide valuable data sources for this purpose6, 49.
+While previous agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways50, 36, 41, 9, there is a growing need to integrate real-world data to strengthen the connections between simulation results and real-world outcomes.
+Social media can provide valuable data sources for this purpose6, 51.
 We address this challenge by proposing a data-informed epidemic model that takes both the distribution of misinformed individuals and a physical mobility network into account.
-Using this data, we augment the Susceptible-Infected-Recovered (SIR) model to account for a subpopulation of “misinformed” individuals, who do not heed expert public health guidance.
+Using this data, we augment the Susceptible-Infected-Recovered (SIR) model to account for a subpopulation of “misinformed” individuals.
 We explore how this group can affect the larger, ordinary population using both a mean-field approximation, which assumes all individuals have an equal chance of interacting, and a multi-level agent-based simulation based on two large, data-informed network: a social network where misinformation spreads and a contact network where the disease can propagate.
@@
 The fractions of individuals in the susceptible, infected, and recovered compartments are denoted by $S_{O}$, $I_{O}$, and $R_{O}$ respectively for the ordinary group; and by $S_{M}$, $I_{M}$, and $R_{M}$ for the misinformed group.
-Ordinary individuals are considered to be well-informed about public health guidelines, while misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow recommended behaviors, such as social distancing, mask-wearing, and vaccination, thereby increasing the risk of infection for themselves and others19.
+Ordinary individuals are considered to be well-informed about public health guidelines, while misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow recommended behaviors, such as social distancing, mask-wearing, and vaccination, thereby increasing the risk of infection for themselves and others20.
 We therefore define two distinct parameters $\beta_{O}$ and $\beta_{M}$ for ordinary and misinformed susceptibles, respectively.
@@
 For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig.
-3C 35.
+3C 37.
 Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig.
@@
 3C): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals.
-In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 44, 47.
+In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 46, 49.
 Agent-based simulations
@@
 (C) The infection spreads through the contact network (black nodes), according to the SMIR model.
-We leverage a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines21.
+We leverage a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines22.
 From approximately nine months of this data (Jan.
@@
 With this data, we build a directed and weighted information diffusion network, in which an edge $(i\rightarrow j,w)$ indicates that $j$ retweeted $i$ $w$ times.
-There are various ways to model the infodemic18.
+There are various ways to model the infodemic19.
 We simulate the spread of misinformation on this network, as illustrated in Fig.
@@
 Not all users actively participate in content sharing, with only about half of U.S.
-Twitter users, for instance, engaging in this activity36.
-But even without active sharing, exposure to misinformation can still influence individual behavior31.
-To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process25.
-This is a simple way to model complex contagion15, 51, 33, according to which individuals require multiple exposures to misinformation before they become misinformed themselves.
+Twitter users, for instance, engaging in this activity38.
+But even without active sharing, exposure to misinformation can still influence individual behavior33.
+To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process27.
+While many social influence models have been proposed14, this is a simple way to capture complex contagion, according to which individuals require multiple exposures to misinformation before they become misinformed themselves16, 54, 35.
 Let a linear threshold $\phi$ represent the minimum number of misinformed friends needed for an ordinary node to become misinformed.
@@
 The process begins by selecting a sample of individuals from each county within the information network.
-As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic27, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S.
+As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic29, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S.
 presidential election.
@@
 Mitigation measures such as social distancing decrease $\bar{k}$, while those such as masking and vaccination decrease $p$.
-A simple way to model the combined effects of misinformation on these behaviors is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic30, and explore a wide range of transmission probabilities to capture worst- and best-case scenarios.
+A simple way to model the combined effects of misinformation on these behaviors is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic32, and explore a wide range of transmission probabilities to capture worst- and best-case scenarios.
 To this end, we model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals.
@@
 This work presents large-scale simulations based on an epidemic model that accounts for a subpopulation of misinformed individuals, informed by real-world data.
-Such an approach answers calls from the literature to anchor models in empirical data6, 49, allowing us to explore realistic potential outcomes.
-While we know that exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal37, our model further links these behaviors by a misinformed population to an accelerated spread of disease.
+Such an approach answers calls from the literature to anchor models in empirical data6, 51, allowing us to explore realistic potential outcomes.
+While we know that exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal39, our model further links these behaviors by a misinformed population to an accelerated spread of disease.
 By deriving the mean-field system of equations for the novel SMIR model and exploring its dynamics, we demonstrate how the risky behaviors of misinformed individuals can adversely impact those following public health guidelines, worsening outcomes for the entire population.
@@
 This would result in an additional 14% of the population becoming infected — nearly 47 million Americans based on recent U.S.
-Census data50.
+Census data52.
 The corresponding price tag of vaccine misinformation would be over \$143B, using estimated health care costs associated with COVID-19 in the U.S.3.
-These results provide an alarming quantitative bound on the harm of exposure to online vaccine misinformation.
+While these figures are based on extreme scenarios, they represent an alarming bound on the harm of exposure to online vaccine misinformation.
 They should provide public health authorities as well as social media platforms with heightened motivation to curb vaccine misinformation, despite the difficulties posed by social media design11.
@@
 Such integration would enable the transition of individuals from ordinary to misinformed susceptible states throughout the simulation, allowing for a simultaneous examination of opinion and disease dynamics.
-Some theoretical models have already explored similar approaches and obtained results that align with our findings34, 48.
+Some theoretical models have already explored similar approaches and obtained results that align with our findings36, 50.
+Future work should attempt to analyze the potential effects of misinformation in more realistic scenarios, where values for the key parameters $p_{M}$ and $p_{O}$ could be obtained by calibrating the model with empirical epidemic data from particular regions and time periods.
 Additional extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population.
@@
 The mean-field model is governed by the following system of equations:
-$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-\beta_{O}S_{O}(I_{O}+I_{M}),\;\;\frac{dI_{O}}{dt}=\beta_{O}S_{O}(I_{O}+I_{M})-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\\ \frac{dS_{M}}{dt}=-\beta_{M}S_{M}(I_{O}+I_{M}),\;\;\frac{dI_{M}}{dt}=\beta_{M}S_{M}(I_{O}+I_{M})-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (1)
+$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-\beta_{O}S_{O}(I_{O}+I_{M}),\;\;\frac{dI_{O}}{dt}=\beta_{O}S_{O}(I_{O}+I_{M})-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{dS_{M}}{dt}=-\beta_{M}S_{M}(I_{O}+I_{M}),\;\;\frac{dI_{M}}{dt}=\beta_{M}S_{M}(I_{O}+I_{M})-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (1)
 To account for homophily, we modify the term $I_{O}+I_{M}$ in Eq.
@@
 We thus obtain:
-$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-2\beta_{O}S_{O}(I_{O}\alpha+I_{M}(1-\alpha)),\;\;\frac{dI_{O}}{dt}=2\beta_{O}S_{O}(I_{O}\alpha+I_{M}(1-\alpha))-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\\ \frac{dS_{M}}{dt}=-2\beta_{M}S_{M}(I_{O}(1-\alpha)+I_{M}\alpha),\;\;\frac{dI_{M}}{dt}=2\beta_{M}S_{M}(I_{O}(1-\alpha)+I_{M}\alpha)-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (2)
+[…]
 We initialize the simulation by randomly marking $\frac{\epsilon}{2}$ nodes as infected in both the misinformed and ordinary groups, where $\epsilon=0.001$.
 Twitter and derived data.
-Twitter posts in the CoVaxxy dataset21 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API).
+Twitter posts in the CoVaxxy dataset22 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API).
 To capture the online discourse surrounding COVID-19 vaccines in English, a comprehensive set of English-language keywords was carefully curated.
-Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique22 was used to identify co-occurring relevant keywords in December 202021.
-The resulting list contained almost 80 keywords, available online20.
+Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique23 was used to identify co-occurring relevant keywords in December 202022.
+The resulting list contained almost 80 keywords, available online21.
 To confirm the relevance of the collected tweets to the topic of vaccines, we examined the coverage obtained by incrementally adding keywords, starting with the most common ones.
-Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library23 that leverages self-reported location metadata within user profiles (embedded in tweets).
+Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library24 that leverages self-reported location metadata within user profiles (embedded in tweets).
 As an account’s location may change over time (captured across multiple tweets), we utilize the most recent location.
@@
 The information network is constructed from accounts in 341 counties that contain more than 200 Twitter users each.
-Political alignment is estimated using a third-party list of annotated news sources42, 41.
+Political alignment is estimated using a third-party list of annotated news sources44, 43.
 It is averaged across all the sources shared by each account.
@@
 Tweets containing links to articles from a list of low-credibility sources compiled by NewsGuard (score below 60) are labeled as spreading misinformation.
-This approach is common practice and has been validated in the literature37, 26, 8, 29, 46.
+This approach is common practice and has been validated in the literature39, 28, 8, 31, 48.
 Contact network edges.
-To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data52, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States.
+To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data55, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States.
 This data has been widely employed to study human mobility patterns during the COVID pandemic.
@@
 We then normalized $L_{xy}$ to obtain the average probability of individuals in counties $x$ and $y$ coming into contact, and multiplied by the total number of edges to obtain the expected number of connections between individuals in counties $x$ and $y$: $E_{xy}=\frac{L_{xy}}{\sum_{x^{\prime},y^{\prime}}L_{x^{\prime}y^{\prime}}}\frac{\bar{k}N}{2}$ where the sum is over all county pairs and $\frac{\bar{k}N}{2}$ is the total number of edges.
-Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model28 used to generate networks with localized communities.
+Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model30 used to generate networks with localized communities.
 For each pair of distinct locations $x$ and $y$, we draw $E_{xy}$ edges between random pairs of nodes in $x$ and $y$.
@@
 The disease spreading dynamics are then simulated for 100 steps, which correspond to days.
-To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period14 ($\gamma=0.2$).
+To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period15 ($\gamma=0.2$).
 Each simulation is repeated ten times, and the average outcome is reported.
@@
 2022BAXSPY), Knight Foundation, and Craig Newmark Philanthropies.
+Author contributions
+The initial concept for this research was developed by YA, SF, AF, FM.
+The final study design was developed collectively by all authors.
+Data collection was led by MRD as part of the CoVaxxy project 21.
+All mean-field modeling was conducted by MRD.
+FP developed the early code for the agent-based simulation, which was subsequently refined and expanded by MRD for publication.
+Analysis was conducted primarily by MRD, with help from FP, and guidance from YA, SF, AF, FM.
+The SMIR model was developed collaboratively by all authors.
+Visualizations were created by MRD, FP, and FM with input from all authors.
+The first draft was written by MRD, with revisions from all authors.
+FM oversaw the progression of the study.
+Competing interests
+Authors declare that they have no competing interests.
 Ethics
@@
 Increasing $\tau$ means that individuals remain infected longer, so the population gets infected faster and the peak infection is reached more rapidly.
-We set $\tau=5$ ($\gamma=0.2$) in the main text to align with quarantine recommendations from the CDC14.
+We set $\tau=5$ ($\gamma=0.2$) in the main text to align with quarantine recommendations from the CDC15.
 With this value and $\beta_{O}=0.3$, the dynamics are at epidemic levels ($R_{0}=1.5$).
@@
 S4 for $5\leq\bar{k}\leq 25$.
-We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions30.
+We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions32.
 As expected, Fig.

### v4

@@
-Modeling the amplification of epidemic spread by misinformed populations
+Modeling the amplification of epidemic spread by individuals exposed to misinformation on social media
 Abstract
@@
 Here, we employ an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data.
-The model allows us to simulate and estimate various scenarios to understand the impact of misinformation on epidemic spreading.
-Using this model, we present a worst-case scenario in which a heavily misinformed population would result in an additional 14% of the U.S.
-population becoming infected over the course of the COVID-19 epidemic, compared to a best-case scenario.
+The model allows us to simulate various scenarios to understand how epidemic spreading can be affected by misinformation spreading through one particular social media platform.
+Using this model, we compare a worst-case scenario, in which individuals become misinformed after a single exposure to low-credibility content, to a best-case scenario where the population is highly resilient to misinformation.
+We estimate the additional portion of the U.S.
+population that would become infected over the course of the COVID-19 epidemic in the worst-case scenario.
+This work can provide policymakers with insights about the potential harms of exposure to online vaccine misinformation.
 Social factors, such as information sharing, play a crucial role in shaping the dynamics and epidemiology of infectious diseases12, 5.
-For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures34, 4.
-A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic40, or by social media and the anti-vaccination movement1, 26, 10, 13.
-A great deal of work has explored how to model the influence of human behaviour on the spread of infectious diseases53, 25.
+For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures36, 4.
+A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic42, or by social media and the anti-vaccination movement1, 27, 10, 14.
+A great deal of work has explored how to model the influence of human behavior on the spread of infectious diseases59, 26.
 Here we focus on risky behaviors affecting disease transmission that are associated with misinformed individuals.
-Misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance45.
-Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States42, 39.
-Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments33.
-This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity17, 47.
+Misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance48.
+Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States45, 41.
+Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments35.
+This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity18, 50.
 Proper management of epidemic crises in the modern age thus requires the understanding of the complex relationship between the spread of (mis)information through online social networks and the spread of disease through physical contact networks (Fig.
@@
 In a scenario where the misinformation spreads more widely (purple nodes in E), more individuals get infected (yellow nodes in F).
-While previous agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways50, 36, 41, 9, there is a growing need to integrate real-world data to strengthen the connections between simulation results and real-world outcomes.
-Social media can provide valuable data sources for this purpose6, 51.
+Agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways54, 38, 44, 9.
+One model estimated that between March and November 2021, misinformation caused at least 198 thousand additional COVID-19 cases, 2,800 additional deaths, and \$299M in additional hospital costs in Canada30.
+However, there is a growing need to strengthen the connections between simulation results and real-world outcomes by integrating real-world data from social media6, 55.
 We address this challenge by proposing a data-informed epidemic model that takes both the distribution of misinformed individuals and a physical mobility network into account.
-Using this data, we augment the Susceptible-Infected-Recovered (SIR) model to account for a subpopulation of “misinformed” individuals.
-We explore how this group can affect the larger, ordinary population using both a mean-field approximation, which assumes all individuals have an equal chance of interacting, and a multi-level agent-based simulation based on two large, data-informed network: a social network where misinformation spreads and a contact network where the disease can propagate.
+Using this data, we augment the Susceptible Infected Recovered (SIR) model to account for a subpopulation of “misinformed” individuals.
+We refer to this as the Susceptible Misinformed Infected Recovered (SMIR) model.
+We explore how the misinformed group can affect the larger, ordinary population using a multi-level agent-based simulation based on two large, data-informed networks: a social network where misinformation spreads and a contact network where the disease can propagate.
 A contact network of approximately 20 million nodes is constructed by leveraging large-scale Twitter data, county-level voting records, and cell phone mobility data.
 We incorporate theoretically extreme values of the parameter responsible for the epidemic transmission to evaluate best- and worst-case scenarios about the impact of misinformed individuals on the spread of disease and obtain quantitative bounds on the harm caused by misinformation.
+The proposed model lets us move beyond simplified experimental settings to assess the impacts of misinformation56.
 Results
-In our model, as in the standard SIR modeling framework2, a parameter $\beta$ describes the average number of infected individuals generated by an infected individual in a time unit.
-Infected individuals recover with rate $\gamma$.
-We extend this epidemic model to account for misinformed and ordinary subpopulations.
-We utilize six compartments to capture the three stages of SIR for each subpopulation.
-The fractions of individuals in the susceptible, infected, and recovered compartments are denoted by $S_{O}$, $I_{O}$, and $R_{O}$ respectively for the ordinary group; and by $S_{M}$, $I_{M}$, and $R_{M}$ for the misinformed group.
-Ordinary individuals are considered to be well-informed about public health guidelines, while misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow recommended behaviors, such as social distancing, mask-wearing, and vaccination, thereby increasing the risk of infection for themselves and others20.
-We therefore define two distinct parameters $\beta_{O}$ and $\beta_{M}$ for ordinary and misinformed susceptibles, respectively.
-We model the risky behavior of the misinformed group by setting $\beta_{M}\geq\beta_{O}$.
-The recovery rate $\gamma$ is identical for both subpopulations.
-We refer to this as the Susceptible Misinformed Infected Recovered (SMIR) model.
-Mean-field simulations
-The mean-field approach assumes that an infected individual can transmit disease to any susceptible individual.
-With this assumption, we derive the model’s mean-field system of equations (see Methods) and explore parameter values for which epidemic spread occurs.
-We denote the proportions of ordinary and misinformed individuals as $\mu$ and $1-\mu$, respectively.
-An initial proportion of the population $\epsilon$ is infected, split evenly between the ordinary and misinformed groups.
-Thus, we initialize the compartments with $I_{O}=\frac{\epsilon}{2}$, $I_{M}=\frac{\epsilon}{2}$, $S_{O}=\mu-\frac{\epsilon}{2}$, $S_{M}=1-\mu-\frac{\epsilon}{2}$, and $R_{O}=R_{M}=0$.
-The effects of varying the transmission and recovery parameters are predictable (see Supplementary Information): lower $\beta$ and higher $\gamma$ “flatten the curve” and reduce the negative outcomes of an infection.
-Based on these explorations we set $\beta_{O}=0.3$ and $\gamma=0.2$ such that the basic reproduction number $R_{0}=\beta/\gamma>1$ to ensure epidemic spread.
-Next we focus on the effects of varying the misinformed subpopulation’s level of risk and the degree of homophily within the network.
-Figure 2: Increasing $\lambda=\beta_{M}/\beta_{O}$ accelerates and amplifies the infection.
-We use $\beta_{O}=0.3$, $\gamma=0.2$, and $\mu=0.5$.
-(A) Proportion of the population infected on each day.
-(B) Extra proportion of the total population that is infected as a function of $\lambda$.
-To explore the effects of risky behaviors by misinformed individuals, let us assume two equally-sized subpopulations ($\mu=1/2$) and introduce the scaling factor $\lambda=\beta_{M}/\beta_{O}\geq 1$.
-Fig.
-2 illustrates the increasing negative impact of the misinformed subpopulation on the disease-spreading dynamics as $\lambda$ becomes larger.
-The social cost associated with the more risky behaviors by the misinformed group is passed on to the whole network.
-For example, when $\lambda=3.0$, peak infection for the entire population is reached 38 days earlier than in the $\lambda=1$ case (day 22 vs.
-60; Fig.
-2A), leading to an additional 29.2% of the population becoming infected (Fig.
-2B).
-Let us explore the effect of homophily among the ordinary and misinformed subpopulation networks.
-Homophily means that infected individuals are more likely to interact with (and infect) susceptibles from the same subpopulation (ordinary or misinformed) than the other group.
-The degree of homophily is modeled by a parameter $\alpha$.
-When $\alpha=0.5$, individuals are equally likely to interact within and across groups (no homophily), whereas $\alpha=1$ is the case when homophily is strongest and the subpopulations do not interact with each other (see Methods for details).
-Figure 3: Homophily in the contact network worsens the infection among misinformed individuals, especially for lower transmission rates.
-The combined effects of transmission and homophily parameters, $\beta_{O}$ and $\alpha$, are examined with the mean-field approximation when $\lambda=3$, $\gamma=0.2$, and $\mu=0.5$.
-We plot the proportions of infected individuals in (A) the ordinary population, (B) the misinformed population, and (C) the overall population.
-The maximum proportion of the overall population infected for each $\beta_{O}$ is marked with a black dot.
-When the transmission rate is sufficiently high, homophily benefits the entire population but harms the misinformed group.
-Fig.
-3 illustrates the effects of homophily ($0.5\leq\alpha\leq 1$) for different levels of ordinary transmission ($0.1\leq\beta_{O}\leq 0.4$).
-For less infectious disease (low $\beta_{O}$), increasing homophily significantly harms the misinformed group (Fig.
-3B): the infection remains confined within this group.
-There is no discernible effect on the ordinary population as long as the two groups interact; when they do not ($\alpha=1$), we observe a sharp reduction in infections (Fig.
-3A).
-For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig.
-3C 37.
-Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig.
-3B and C).
-As $\beta_{O}$ increases further, while nearly the entire misinformed population becomes infected regardless of $\alpha$ (Fig.
-3B), high homophily shields the full population (Fig.
-3C): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals.
-In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 46, 49.
-Agent-based simulations
-To examine the influence of misinformation within a more realistic epidemic framework, we utilize a multi-level model.
+We utilize a multi-level, agent-based model to examine the influence of misinformation on epidemic spread.
 Our approach combines an empirically derived information network with a contact network calibrated with real-world data, as illustrated in Fig.
-4.
-Information diffusion is modeled by leveraging a large set of users of a popular social media platform and epidemic simulations are subsequently conducted on contact networks populated with misinformed individuals.
+2.
+Information diffusion is modeled by leveraging a large set of users of a popular social media platform.
+Epidemic simulations are subsequently conducted on contact networks populated with misinformed individuals.
 Figure 2: An idealized example of our multi-level modeling framework.
@@
 (C) The infection spreads through the contact network (black nodes), according to the SMIR model.
-We leverage a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines22.
+We start from a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines23.
 From approximately nine months of this data (Jan.
@@
 population, and people also access information in other ways, such as traditional media and word of mouth.
-However, this empirical social network serves as a data-informed proxy for a large, realistic network through which people share information about the disease.
+However, this social media platform serves as one large, realistic network through which people share information about the disease.
 With this data, we build a directed and weighted information diffusion network, in which an edge $(i\rightarrow j,w)$ indicates that $j$ retweeted $i$ $w$ times.
-There are various ways to model the infodemic19.
+There are various ways to model the infodemic20.
 We simulate the spread of misinformation on this network, as illustrated in Fig.
-4A.
+2A.
 Accounts that share or reshare posts containing misinformation are considered misinformed.
-These accounts function as the initial seeds from which misinformation proliferates to the wider network.
-Not all users actively participate in content sharing, with only about half of U.S.
-Twitter users, for instance, engaging in this activity38.
-But even without active sharing, exposure to misinformation can still influence individual behavior33.
-To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process27.
-While many social influence models have been proposed14, this is a simple way to capture complex contagion, according to which individuals require multiple exposures to misinformation before they become misinformed themselves16, 54, 35.
+These accounts serve as the initial seeds from which misinformation proliferates, with exposure to this content likely concentrated within the wider network13.
+Many users may not actively participate in content sharing; for instance, only about half of U.S.
+Twitter users engage in sharing40.
+Even without active sharing, exposure to misinformation or misleading content can still influence individual behavior35, 1.
+To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process28.
+While many social influence models have been proposed15, this is a simple way to capture complex contagion, according to which individuals may require multiple exposures to misinformation before they become misinformed themselves17, 60, 37.
 Let a linear threshold $\phi$ represent the minimum number of misinformed friends needed for an ordinary node to become misinformed.
@@
 We can interpret $\phi$ as a measure of “resilience” to misinformation; as it increases, individuals require more exposure to misinformation to be converted to the misinformed group.
+Conversely, we can think of $\phi$ as inversely related to intent or motivation to engage with low-credibility content52.
 Note that since we explore the full range of $\phi$ values, the following results are unaffected whether the threshold is defined based on the number of users or the number of retweets.
 Fig.
-5A shows how $\phi$ influences the number of misinformed individuals within the retweet network.
+3A shows how $\phi$ influences the number of misinformed individuals within the retweet network.
 With strong resilience ($\phi>10$), exposure to misinformation does not have much effect and few nodes are converted to the misinformed group.
@@
 Given a threshold $\phi$ and the corresponding information network, we aim to construct a physical contact network containing empirically calibrated misinformed subpopulations (Fig.
-4B).
+2B).
 The process begins by selecting a sample of individuals from each county within the information network.
-As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic29, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S.
+As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic31, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S.
 presidential election.
@@
 This process captures empirical measurements of the ideological split, relative population size, and quantity of misinformed individuals in each county.
+It also allows us to account for the known link between the ideological motivations of users and their exposure to misinformation13, 41.
 We add contact network edges by leveraging cell phone mobility data that provides the probability of an individual traveling within and between counties.
 See Methods for details.
-Disease-spreading dynamics on a contact network are simulated using an agent-based version of SMIR (Fig.
-4C).
-Two critical parameters affecting these dynamics are the density of the contact network, captured by its average degree $\bar{k}$, and the transmission probability $p$ (note that $p\bar{k}=\beta$).
+Disease-spreading dynamics on the contact network are simulated using the SMIR model (Fig.
+2C).
+As in the standard SIR2, a parameter $\beta$ describes the average number of infected individuals generated by an infected individual in a time unit.
+We can express $\beta=p\bar{k}$ in terms of two critical parameters that affect the spreading dynamics: the density of the contact network, captured by its average degree $\bar{k}$, and the transmission probability $p$.
+Infected individuals recover with rate $\gamma$.
+We extend this epidemic model to account for misinformed and ordinary subpopulations.
+Ordinary individuals are considered to be well-informed about public health guidelines, such as social distancing, mask-wearing, and vaccination.
 Mitigation measures such as social distancing decrease $\bar{k}$, while those such as masking and vaccination decrease $p$.
-A simple way to model the combined effects of misinformation on these behaviors is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic32, and explore a wide range of transmission probabilities to capture worst- and best-case scenarios.
-To this end, we model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals.
+Misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow these recommended behaviors, thereby increasing the risk of infection for themselves and others21.
+A simple way to model the combined effects of misinformation on these behaviors through a single parameter is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic34, and use extreme values of $p$ to capture worst- and best-case scenarios.
+An effective reduction of contacts, resulting for example from social distancing or lockdowns, can be captured by decreasing the $p$ parameter.
+We therefore model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals.
 In contrast, we model the adoption of several mitigation measures by selecting an extremely small value $p_{O}=0.01$ for ordinary individuals.
-These values illustrate two theoretically extreme scenarios.
-The former portrays a realistic number of interactions during non-pandemic times, accompanied by high transmission rates due to the absence of preventive measures, such as social distancing, mask-wearing, or vaccinations.
+The former scenario portrays a realistic number of interactions during non-pandemic times, accompanied by high transmission rates due to the absence of preventive measures, such as social distancing, mask-wearing, or vaccinations.
 The latter demonstrates decreased daily interactions and reduced transmission rates resulting from the implementation of these preventive measures.
 Using the empirically calibrated contact networks in conjunction with these extreme parameters, the simulation approach allows us to bound the best- and worst-case scenarios in a data-informed manner (see Methods for more information).
-The worst-case effect of the misinformed subpopulation size on the daily incidence of infection (illustrated in Fig.
-1C-F on a small network) is quantified in Fig.
-5B on a large network (10% sample).
-A heavily misinformed population ($\phi=1$) corresponds to an additional 9% of the population being infected at peak time (a six-fold increase) compared to a population following expert guidance in the best-case scenario ($\phi=20$).
+The effects of the misinformed subpopulation size on the daily incidence of infection (illustrated in Fig.
+1C-F on a small network) are quantified in Fig.
+3B on a large network (10% sample).
+The worst case capturing a heavily misinformed population ($\phi=1$) corresponds to an additional 9% of the population being infected at peak time (a six-fold increase) compared to a resilient population following expert guidance in the best-case scenario ($\phi=20$).
 The peak also occurs approximately two weeks earlier.
-The cumulative effect of a heavily misinformed population is also significant, with an additional 14% of the population infected over the course of the epidemic — a 32% relative increase (Fig.
-5C).
-We repeated this analysis for different sample sizes and found that the main results are robust (see Supplementary Information).
+The cumulative effect is also significant, with an additional 14% of the population infected over the course of the epidemic compared to case with a more resilient population — a 32% relative increase (Fig.
+3C).
+We explored alternative scenarios for the ratio $p_{M}/p_{O}$ through a mean-field approximation.
+Predictably, as this ratio gets larger, the infected population increases and the peak infection occurs earlier.
+We also considered different sample sizes for the empirical network and found that the main results are robust.
+These analyses can be found in Supplementary Information.
 Discussion
-This work presents large-scale simulations based on an epidemic model that accounts for a subpopulation of misinformed individuals, informed by real-world data.
-Such an approach answers calls from the literature to anchor models in empirical data6, 51, allowing us to explore realistic potential outcomes.
-While we know that exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal39, our model further links these behaviors by a misinformed population to an accelerated spread of disease.
-By deriving the mean-field system of equations for the novel SMIR model and exploring its dynamics, we demonstrate how the risky behaviors of misinformed individuals can adversely impact those following public health guidelines, worsening outcomes for the entire population.
-Additionally, increasing homophily can benefit the overall population by protecting ordinary citizens; however, it may also lead to higher infection rates within the misinformed subpopulation.
-Agent-based simulations of the SMIR model let us study the epidemic on empirically calibrated contact networks, which provide more realistic scenarios compared to the mean-field approach.
-The model estimates that, in the worst-case scenario, a heavily misinformed population can amplify the peak of the infection by a factor of six and accelerate it by two weeks.
+Exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal41.
+There is also experimental evidence suggesting a causal link35, 57, 1.
+While one study found no evidence that misinformation reduces intent to vaccinate, the authors report that they did not have sufficient power to detect small effects43.
+Assuming an association exists between exposure to health misinformation on one particular social media platform and risky behaviors, this work uses large-scale epidemic simulations to further link the behaviors of misinformed individuals to an accelerated spread of disease.
+Our model is anchored in empirical data6, 55 to explore potential outcomes.
+Agent-based simulations of the SMIR model let us study the epidemic on empirically calibrated contact networks.
+By comparing a worst-case scenario, in which individuals become misinformed after a single exposure to low-credibility content, to a best-case scenario where the population is highly resilient to misinformation, the model estimates that the peak of the infection is amplified by a factor of six and accelerated by two weeks.
 This would result in an additional 14% of the population becoming infected — nearly 47 million Americans based on recent U.S.
-Census data52.
+Census data58.
 The corresponding price tag of vaccine misinformation would be over \$143B, using estimated health care costs associated with COVID-19 in the U.S.3.
@@
 They should provide public health authorities as well as social media platforms with heightened motivation to curb vaccine misinformation, despite the difficulties posed by social media design11.
-We acknowledge certain limitations in our approach.
-While individual beliefs and behaviors may vary over time, our model simplifies the scenario by dichotomizing individuals into misinformed and ordinary subpopulations and assuming constant transmission rates.
+Our results do not address the differential effects of the epidemic on the two populations of ordinary and misinformed individuals.
+We carry out such an analysis using a mean-field approximation of the model, which assumes all individuals have an equal chance of interacting.
+The mean-field model demonstrates how the risky behaviors of misinformed individuals can adversely impact those following public health guidelines, worsening outcomes for the entire population (see Supplementary Information).
+Additionally, we use the mean-field model to explore the role of homophily in the population, i.e., scenarios where misinformed individuals are more likely to be connected to other misinformed individuals and similarly for the ordinary population.
+We find that increasing homophily can benefit the overall population by protecting ordinary citizens; however, it may also lead to higher infection rates within the misinformed subpopulation (see Supplementary Information).
+We acknowledge several limitations in our approach.
+The model assumes the existence of a causal link between exposure to online misinformation and the adoption of risky behaviors.
+There is a need for models that can provide support for this assumption beyond existing lab experiments35, 1.
+Using empirical retweet data as a proxy for social connections may not capture potential passive exposure to misinformation.
+While follower relationships could diminish this limitation, our choice allows us to focus on users who are more likely to be impacted due to their active engagement.
+We model a single wave of infection with somewhat arbitrary extreme-case parameters ($p_{O}=0.01,p_{M}=1$).
+A broader range of values is explored in a mean-field scenario, along with the effect of the size of the misinformed population (see Supplementary Information).
+Of course, as $p_{O}\rightarrow 0$, only the misinformed population can get infected in the model.
+However, since the mean-field scenario ignores the network structure, its results cannot be directly compared to those of the agent-based model.
+COVID-19 saw multiple waves of infection with different variants, varying reproduction numbers, levels of immunity, and so on.
+Future work should attempt to quantify the potential effects of misinformation in more realistic scenarios, where the key parameters $p_{M}$ and $p_{O}$ could be calibrated on empirical surveillance data from particular regions and time periods.
 We also assume uniform resilience to misinformation for all individuals during the information diffusion process, although this attribute likely differs across individuals.
@@
 Such integration would enable the transition of individuals from ordinary to misinformed susceptible states throughout the simulation, allowing for a simultaneous examination of opinion and disease dynamics.
-Some theoretical models have already explored similar approaches and obtained results that align with our findings36, 50.
-Future work should attempt to analyze the potential effects of misinformation in more realistic scenarios, where values for the key parameters $p_{M}$ and $p_{O}$ could be obtained by calibrating the model with empirical epidemic data from particular regions and time periods.
-Additional extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population.
+Some theoretical models have already explored similar approaches and obtained results that align with our findings38, 54.
+Finally, although individual beliefs and behaviors may vary over time, our model simplifies the scenario by dichotomizing individuals into misinformed and ordinary subpopulations and assuming constant transmission rates.
+Future extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population62.
 Methods
-SMIR model.
-For both the ordinary and misinformed subpopulations, the Susceptible Misinformed Infected Recovered (SMIR) model replicates the standard SIR compartments, denoted as $S_{O}$/$I_{O}$/$R_{O}$ and $S_{M}$/$I_{M}$/$R_{M}$, respectively.
-While SMIR adopts distinct transmission parameters for the misinformed and ordinary groups ($\beta_{M}$ and $\beta_{O}$, respectively), the mean-field approach assumes that infected individuals from either group ($I_{O}+I_{M}$) can potentially infect anyone in the susceptible populations.
-The mean-field model is governed by the following system of equations:
-$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-\beta_{O}S_{O}(I_{O}+I_{M}),\;\;\frac{dI_{O}}{dt}=\beta_{O}S_{O}(I_{O}+I_{M})-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{dS_{M}}{dt}=-\beta_{M}S_{M}(I_{O}+I_{M}),\;\;\frac{dI_{M}}{dt}=\beta_{M}S_{M}(I_{O}+I_{M})-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (1)
-To account for homophily, we modify the term $I_{O}+I_{M}$ in Eq.
-1 to account for increased (decreased) contacts within (across) groups, according to the parameter $\alpha\geq 0.5$.
-When homophily does not play a role ($\alpha=0.5$), there is an equal probability of interacting with either subpopulation’s infected group.
-We thus obtain:
-[…]
-We initialize the simulation by randomly marking $\frac{\epsilon}{2}$ nodes as infected in both the misinformed and ordinary groups, where $\epsilon=0.001$.
 Twitter and derived data.
-Twitter posts in the CoVaxxy dataset22 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API).
+Twitter posts in the CoVaxxy dataset23 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API).
 To capture the online discourse surrounding COVID-19 vaccines in English, a comprehensive set of English-language keywords was carefully curated.
-Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique23 was used to identify co-occurring relevant keywords in December 202022.
-The resulting list contained almost 80 keywords, available online21.
+Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique24 was used to identify co-occurring relevant keywords in December 202023.
+The resulting list contained almost 80 keywords, available online22.
 To confirm the relevance of the collected tweets to the topic of vaccines, we examined the coverage obtained by incrementally adding keywords, starting with the most common ones.
-Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library24 that leverages self-reported location metadata within user profiles (embedded in tweets).
+Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library25 that leverages self-reported location metadata within user profiles (embedded in tweets).
 As an account’s location may change over time (captured across multiple tweets), we utilize the most recent location.
@@
 The information network is constructed from accounts in 341 counties that contain more than 200 Twitter users each.
-Political alignment is estimated using a third-party list of annotated news sources44, 43.
+Political alignment is estimated using a third-party list of annotated news sources47, 46.
 It is averaged across all the sources shared by each account.
@@
 Tweets containing links to articles from a list of low-credibility sources compiled by NewsGuard (score below 60) are labeled as spreading misinformation.
-This approach is common practice and has been validated in the literature39, 28, 8, 31, 48.
+This approach is common practice and has been validated in the literature41, 29, 8, 33, 51.
 Contact network edges.
-To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data55, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States.
+To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data61, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States.
 This data has been widely employed to study human mobility patterns during the COVID pandemic.
@@
 We then normalized $L_{xy}$ to obtain the average probability of individuals in counties $x$ and $y$ coming into contact, and multiplied by the total number of edges to obtain the expected number of connections between individuals in counties $x$ and $y$: $E_{xy}=\frac{L_{xy}}{\sum_{x^{\prime},y^{\prime}}L_{x^{\prime}y^{\prime}}}\frac{\bar{k}N}{2}$ where the sum is over all county pairs and $\frac{\bar{k}N}{2}$ is the total number of edges.
-Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model30 used to generate networks with localized communities.
+Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model32 used to generate networks with localized communities.
 For each pair of distinct locations $x$ and $y$, we draw $E_{xy}$ edges between random pairs of nodes in $x$ and $y$.
@@
 The disease spreading dynamics are then simulated for 100 steps, which correspond to days.
-To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period15 ($\gamma=0.2$).
+To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period16 ($\gamma=0.2$).
 Each simulation is repeated ten times, and the average outcome is reported.
 Data and code availability.
-Code and data are available in a public repository (github.iu.edu/NaN-team/bounding-misinfo-impact-on-disease-spread).
+Code and data are available in a public repository (github.com/osome-iu/bounding-misinfo-impact-on-disease-spread).
 Acknowledgments
@@
 The final study design was developed collectively by all authors.
-Data collection was led by MRD as part of the CoVaxxy project 21.
+Data collection was led by MRD as part of the CoVaxxy project 22.
 All mean-field modeling was conducted by MRD.
@@
 Appendix 1 Mean-field SMIR model
+For both the ordinary and misinformed subpopulations, the Susceptible Misinformed Infected Recovered (SMIR) model replicates the standard SIR compartments, denoted as $S_{O}$/$I_{O}$/$R_{O}$ and $S_{M}$/$I_{M}$/$R_{M}$, respectively.
+SMIR adopts distinct transmission parameters for the misinformed ($\beta_{M}$) and ordinary ($\beta_{O}$) groups.
+(In the agent-based model, these are proportional to $p_{M}$ and $p_{O}$, respectively.) The mean-field approximation assumes that the population is well mixed, ignoring the empirical network structure, and that infected individuals from either group ($I_{O}+I_{M}$) can potentially infect anyone in the susceptible populations.
+The mean-field model is governed by the following system of equations:
+$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-\beta_{O}S_{O}(I_{O}+I_{M}),\;\;\frac{dI_{O}}{dt}=\beta_{O}S_{O}(I_{O}+I_{M})-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{dS_{M}}{dt}=-\beta_{M}S_{M}(I_{O}+I_{M}),\;\;\frac{dI_{M}}{dt}=\beta_{M}S_{M}(I_{O}+I_{M})-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (1)
+To model homophily, we modify the term $I_{O}+I_{M}$ in Eq.
+1 to account for increased (decreased) contacts within (across) groups, according to the parameter $\alpha\geq 0.5$.
+When homophily does not play a role ($\alpha=0.5$), there is an equal probability of interacting with either subpopulation’s infected group.
+We thus obtain:
+[…]
 Let us denote the proportions of misinformed and ordinary individuals as $\mu$ and $1-\mu$, respectively.
@@
 Figure S1: Reducing the transmission parameter $\beta_{O}$ decreases the severity of the epidemic.
-We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values, and (c) the total proportion of the population infected as $\beta_{O}$ varies.
+We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values for the ordinary population, and (c) the total proportion of the population infected as $\beta_{O}$ varies.
+In (a), the curve for $\beta_{O}=0.2$ is difficult to see because the proportion of the population infected remains very low throughout the simulation.
 Here we do not consider the role of misinformation or homophily.
@@
 Figure S2: Effects of varying the recovery rate.
-We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values, and (c) the total proportion of the population infected as a function of the number of days to recover, $\tau$.
+We plot (a) the proportion of the population infected each day, (b) $R_{0}$ values for the ordinary population, and (c) the total proportion of the population infected as a function of the number of days to recover, $\tau$.
 Here we do not consider the role of misinformation or homophily.
@@
 Increasing $\tau$ means that individuals remain infected longer, so the population gets infected faster and the peak infection is reached more rapidly.
-We set $\tau=5$ ($\gamma=0.2$) in the main text to align with quarantine recommendations from the CDC15.
-With this value and $\beta_{O}=0.3$, the dynamics are at epidemic levels ($R_{0}=1.5$).
+In summary, the effects of varying the transmission and recovery parameters are predictable: lower $\beta$ and higher $\gamma$ “flatten the curve” and reduce the negative outcomes of an infection.
+Based on these explorations we set $\tau=5$ ($\gamma=0.2$) to align with quarantine recommendations from the CDC16.
+We further set $\beta_{O}=0.3$ such that the basic reproduction number is $R_{0}\geq\beta_{O}/\gamma=1.5$ to ensure epidemic spread within the ordinary population.
+Appendix 2 Mean-field analyses
+Figure S3: Increasing $\lambda=\beta_{M}/\beta_{O}$ accelerates and amplifies the infection.
+We use $\gamma=0.2$, and $\mu=0.5$.
+(a) Overall proportion of the population infected as a function of $\lambda$, for different values of $\beta_{O}$.
+(b) Proportion of the population infected on each day, for different values of $\lambda$ using $\beta_{O}=0.3$.
+(c) Extra proportion of the total population that is infected as a function of $\lambda$ ($\beta_{O}=0.3$).
+To explore the effects of risky behaviors by misinformed individuals, let us assume two equally-sized subpopulations ($\mu=1/2$) and introduce the scaling factor $\lambda=\beta_{M}/\beta_{O}\geq 1$.
+Fig.
+S3(a) illustrates the increasing negative impact of the misinformed subpopulation on the disease-spreading dynamics as $\lambda$ becomes larger.
+If the ordinary population has very low $\beta_{O}$, $\lambda$ has to be very high for the misinformed population to have an effect.
+On the other hand, if $\beta_{O}$ is large enough, increasing $\lambda$ leads to a ceiling effect, as $\beta_{M}$ cannot exceed one.
+The social cost associated with the more risky behaviors by the misinformed group is passed on to the whole network.
+For example, when $\lambda=3.0$, peak infection for the entire population is reached 38 days earlier than in the $\lambda=1$ case (day 22 vs.
+60; Fig.
+S3(b)), leading to an additional 29.3% of the population becoming infected (Fig.
+S3(c)).
+Figure S4: Increasing the initial proportion $\mu$ of the population that is misinformed, as well as $\beta_{O}$, increases (a) the size of the network that becomes infected and (b) the average $R_{0}$ across the population.
+Here, we fix $\lambda=100$ to match the ratio used in the main text.
+We further explore how the initial size $\mu$ of the misinformed population affects the total proportion of the network that ultimately gets infected.
+We consider various values of $\beta_{M}$ and $\beta_{O}$ such as to capture the same $\lambda=\beta_{M}/\beta_{O}=p_{M}/p_{O}=100$ as in the main text.
+When $\mu$ and $\beta_{O}$ are both low, the misinformed population has no impact on the infection (Fig.
+S4(a)), as $R_{0}<1$ (Fig.
+S4(b)).
+However, increasing either parameter crosses the epidemic threshold ($R_{0}>1$) so that a significant portion of the population gets infected.
+Appendix 3 Effect of homophily
+Let us explore the effect of homophily among the ordinary and misinformed subpopulation networks.
+Homophily means that infected individuals are more likely to interact with (and infect) susceptibles from the same subpopulation (ordinary or misinformed) than the other group.
+The degree of homophily is modeled by a parameter $\alpha$.
+When $\alpha=0.5$, individuals are equally likely to interact within and across groups (no homophily), whereas $\alpha=1$ is the case when homophily is strongest and the subpopulations do not interact with each other (see Methods for details).
+Figure S5: Homophily in the contact network worsens the infection among misinformed individuals, especially for lower transmission rates.
+The combined effects of transmission and homophily parameters, $\beta_{O}$ and $\alpha$, are examined with the mean-field approximation when $\lambda=3$, $\gamma=0.2$, and $\mu=0.5$.
+We plot the proportions of infected individuals in (a) the ordinary population, (b) the misinformed population, and (c) the overall population.
+The maximum proportion of the overall population infected for each $\beta_{O}$ is marked with a black dot.
+When the transmission rate is sufficiently high, homophily benefits the entire population but harms the misinformed group.
+Fig.
+S5 illustrates the effects of homophily ($0.5\leq\alpha\leq 1$) for different levels of ordinary transmission ($0.1\leq\beta_{O}\leq 0.4$).
+For less infectious disease (low $\beta_{O}$), increasing homophily significantly harms the misinformed group (Fig.
+S5(b)): the infection remains confined within this group.
+There is no discernible effect on the ordinary population as long as the two groups interact; when they do not ($\alpha=1$), we observe a sharp reduction in infections (Fig.
+S5(a)).
+For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig.
+S5(c) 39.
+Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig.
+S5B and C).
+As $\beta_{O}$ increases further, while nearly the entire misinformed population becomes infected regardless of $\alpha$ (Fig.
+S5(b)), high homophily shields the full population (Fig.
+S5(c)): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals.
+In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 49, 53.
 Appendix 4 Robustness analyses
 To test the robustness of the main results (Fig.
-5C) with respect to the sample size used to construct the contact network, all simulations were rerun after generating contact networks based on the different sampling percentages between 0.01% and 10%.
+3C) with respect to the sample size used to construct the contact network, all simulations were rerun after generating contact networks based on the different sampling percentages between 0.01% and 10%.
 Fig.
-S3 shows the relative increase in the percentage of the population that becomes infected as a function of the linear threshold $\phi$, using the best-case scenario in which the fewest nodes in the network are misinformed as the baseline.
+S6 shows the relative increase in the percentage of the population that becomes infected as a function of the linear threshold $\phi$, using the best-case scenario in which the fewest nodes in the network are misinformed as the baseline.
 We observe a substantial decrease in the effect of misinformation as the sampling size grows to 1%.
@@
 The $\phi=20$ scenario, in which the fewest nodes in the network are misinformed, is utilized as the baseline.
-We also investigate the impact of the contact network density ($\bar{k}$) on infection dynamics, as illustrated in Fig.
-S4 for $5\leq\bar{k}\leq 25$.
-We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions32.
+Fig.
+S7 illustrates the impact of the contact network density ($\bar{k}$) on infection dynamics, for $5\leq\bar{k}\leq 25$.
+We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions34.
 As expected, Fig.
-S4(a) demonstrates that higher $\bar{k}$ leads to increased infections through the population, since the higher contact density provides more opportunities for transmission.
+S7(a) demonstrates that higher $\bar{k}$ leads to increased infections through the population, since the higher contact density provides more opportunities for transmission.
 But while a larger percentage of the overall population is infected, the relative effect of misinformed individuals decreases.
@@
 The combined effect of these two opposing trends, as shown in Fig.
-S4(b), is that the additional percentage of infected individuals relative to the $\phi=20$ scenario reaches a maximum for some intermediate $\bar{k}$.
+S7(b), is that the additional percentage of infected individuals relative to the $\phi=20$ scenario reaches a maximum for some intermediate $\bar{k}$.
 Fig.
-S4(b) also shows that, consistent with our primary findings, larger $\phi$ (misinformed resilience) results in a smaller infected population.
-In our main analysis we focus on $\bar{k}=25$ and model an effective reduction of contacts, resulting for example from social distancing or lockdowns, by decreasing the $p$ parameter.
+S7(b) also shows that, consistent with our primary findings, increasing $\phi$ (misinformed resilience) decreases the infected population.
+In our main analysis we focus on $\bar{k}=25$ and model an effective reduction of contacts by decreasing the $p$ parameter.
 Figure S7: Effects of average contact network degree $\bar{k}$ on infection dynamics.